What is the wavelength of the photon emitted by a hydrogen atom when a bound electron in the n = 3 state makes the transition to the n = 2 state? (to 2 s.f in nm)
1
Expert's answer
2020-06-01T14:18:34-0400

The wavelength of emitted photon by a hydrogen atom when a bound electron in the n = 3 state makes the transition to the n = 2 state is

"\frac{1}{ \lambda }=R[ \frac{1}{n_1^2}- \frac{1}{n_2^2}]"

"=(1.097*10^7 /m)[ \frac{1}{2^2}- \frac{1}{3^2}]"

"=(1.097*10^7 \/m)[ 0.13889]"

"\lambda=6.56*10^{-7} m=656 nm"

Hence, the wavelength of emitted photon is "\lambda =660 nm"